A 3,000 kilogram car hit with a force of 40,000 Newtons
1
Expert's answer
2020-03-30T07:52:31-0400

The acceleration will be:


"a=\frac{F}{m}=\frac{40000}{3000}=13.3\frac{m}{s^2}"